# ORCID: $0$-0003-2317-2297

### v1

@@ Abstract
Automatic summarisation has been used efficiently in recent years to condense texts, conversations, audio, code, and various other artefacts. A range of methods, from simple template-based summaries to complex machine learning techniques—and more recently, large language models—have been employed to generate these summaries. Summarising software design patterns is important because it helps developers quickly understand and reuse complex design concepts, thereby improving software maintainability and development efficiency. However, the generation of summaries for software design patterns has not yet been explored. Our approach utilises code features and JavaParser to parse the code and create a JSON representation. Using an NLG library on this JSON representation, we convert it into natural language text that acts as a summary of the code, capturing the contextual information of the design pattern. Our empirical results indicate that the summaries generated by our approach capture the context in which patterns are applied in the codebase. Statistical evaluations demonstrate that our summaries closely align with human-written summaries, as evident from high values in the ROUGE-L, BLEU-4, NIST, and FrugalScore metrics. A follow-up survey further shows that DPS summaries were rated as capturing context better than human-generated summaries.
Keywords:
Design Patterns Code Features Summarisation JSON JavaParser
@@ 1 Introduction
  • We are the first to generate design pattern summaries.
  • We propose a novel code feature-based approach that identifies design patterns and constructs summaries containing contextual information.
  • We have built a corpus of approximately 1,500 Java files for experimentation.
  • An evaluation of our approach against human-generated summaries demonstrates statistically significant results, indicating the strong potential of our method.
  • An anonymous survey reveals that our summaries capture design pattern context more effectively than human-generated summaries.
The paper is organised into the following sections. Section 1 introduces the problem, and Section 2 provides the motivation behind the paper, particularly by discussing what context means. Section 3 discusses the experimental setup, focusing on corpus selection and generation, the code features used, the statistical measures for evaluation, and the research questions. Section 4 describes our three-phase summariser approach—DPS. Section 5 evaluates our work based on statistical measures and survey-based feedback. Section 6 discusses threats to internal and external validity, while Section 7 provides a description of related work in design pattern detection and code summarisation. Section 8 concludes our paper and discusses future directions for this research.
@@ 2 Background:Context
McBurney et al. Mcburney:2014 defined context as the purpose of a method within the code in which it exists. This context includes the dependencies of the method and other methods that rely on it. In the context of design patterns, the term refers to the purpose of a class within the pattern—specifically, the role it plays in the class hierarchy, such as whether it acts as a subscriber or a publisher (as in the Observer pattern), as well as the methods invoked within these classes. Although McBurney’s approach focuses solely on the context of the method, we argue that the context of the class is more critical when dealing with design patterns. However, we capture both class and method contexts in our summaries. Thus, we break down the context from the class level to the methods. The following example, taken from the RefactoringGuru GitHub repository, [^2] https://github.com/RefactoringGuru/design-patterns-java/blob/main/src/refactoring_guru/abstract_factory/example/factories/GUIFactory.java accessed and verified on 31-03-2025 illustrates the context.
@@
This code example illustrates that a GUI factory produces two types of products, namely Button and Checkbox, as shown in the implementation class below:
@@
The code presented here illustrates how developers can create families of related objects without specifying their concrete classes. The GUIFactory interface establishes a contract for producing two types of products: Button and Checkbox. Concrete factories, such as MacOSFactory, implement this interface to generate platform-specific variants (MacOSButton and MacOSCheckbox). This approach encapsulates the creation logic, enhancing flexibility and scalability. It not only helps developers identify reusable components, but it also supports refactoring efforts, improves maintainability, and promotes the consistent application of design principles across the codebase. Table 1 shows examples of design pattern contextual summaries of the code examples taken from the corpus.
@@
1 public class EmailNotificationListener implements EventListener { private String email; public EmailNotificationListener(String email) { this.email = email; } @Override public void update(String eventType, File file) { System.out.println("Email to " + email + ": Someone has performed " + eventType + " operation with the following file: " + file.getName()); } } EmailNotificationListener acts as an observer, which updates String parameter of eventType, File parameter of file and sends notifications. It is a public class that implements EventListener. The only method of EmailNotificationListener is update (void). No methods call for update and update calls no methods.
@@
4 public class HtmlDialog extends Dialog { @Override public Button createButton() { return new HtmlButton(); } } HtmlDialog does not have any design pattern. It is a public class that extends Dialog. The only method of HtmlDialog is createButton (Button). No methods call createButton and createButton calls no methods.
5 public class BusAdapter implements Car { private Bus bus; public BusAdapter() { this.bus = new Bus(); } @Override public void drive() { bus.run(); } } BusAdapter acts as an Adapter, which overrides driveCar. It is a public class that implements Car. The only method of BusAdapter is drive (void). No methods call drive and drive calls no methods.
@@ 3.1 Data Collection
To construct our corpus, we manually selected ten open-source GitHub repositories using the keyword “design patterns.” These repositories were not general-purpose software projects but were specifically created to illustrate or teach software design patterns. Most of them originate from educational settings (e.g., university course materials) or accompany books and instructional websites – such as the widely used RefactoringGuru repository. While some repositories contained examples for all 23 Gang of Four (GoF) design patterns, others included a broader or narrower set. We manually curated each repository to retain only those directories relevant to the nine design patterns targeted in our study. This process involved selecting example code files from the appropriate directories and discarding test cases, non-Java files, or unrelated examples. The resulting curated dataset, which we refer to as the DPS-Corpus, contains approximately 1,500 Java files.
Table 2 presents the project names, their URLs, and the number of Java files containing pattern details for the nine design patterns that we aim to investigate. The table summarises the distribution of design pattern implementations across the 10 Java projects within our corpus. Each project demonstrates various design patterns, including Abstract Factory, Adapter, Decorator, Facade, Factory Method, Memento, Observer, Singleton, and Visitor. The total number of pattern files per project ranges from 46 to 943, with the project by shihyu having the highest count at 943 files, and the project by shusheng007 having the lowest count at 46 files. Popular patterns such as Singleton and Factory Method are widely implemented, while others like Memento are less common.
@@ 3.2 Code Features
Inspired by the work on Design Pattern Detection by Nazar et al. Nazar:2022 , where they employ code features to identify design patterns, we have used code features to build the summaries. In this study, we have increased the number of features from 15 to 20 and renamed some of the features used by Nazar et al. While some features (and their names) remain the same, others, such as NOML, are newly introduced. Additionally, features such as class name (CN) and method name (MN) have been renamed to enhance clarity and understanding. These features can be categorised into basic, object-oriented, and binary features, and further subdivided into class-level and method-level features. Feature extraction is a crucial aspect of code analysis, and accurately identifying a design pattern necessitates thorough code analysis. Table LABEL:tab:features lists the selected feature types, feature abbreviations, names, and descriptions.
@@
OO EXtendsFrom (EXF) List of classes that the Java file extends from, i.e. inherits
@@ 3.3.1 BLEU-4
[…]
@@ 3.3.2 NIST
$\text{NIST}=\sum_{n=1}^{N}\left\{\frac{\sum_{\text{all }n\text{-gram match that reference}}\text{Info}(n\text{-gram})}{\sum_{n\text{-gram reference}}}\right\}\cdot\exp(\beta)\log^{2}\left[\min\left(\frac{|\rho|}{|r|},1\right)\right]$ (2)
@@ 3.3.3 ROUGE-L
[…]
@@ 3.3.4 FrugalScore
[…]
@@ 4.1 Parsing
Figure 1: JSON representation of the ComputerPart class from AbdurRKhalid project
@@ 4.3 Summary Generator
We used SimpleNLG [^10] https://github.com/simplenlg/simplenlg, accessed and verified on 23/03/2025, a Java-based natural language generation API, to produce design pattern summaries. This framework leverages lexical resources and realisation components to construct grammatically correct English sentences from structured inputs. The JSON output generated by our Pattern Checker module serves as the input for this phase, with SimpleNLG transforming the structured pattern information into coherent natural language summaries.
@@
Figure 2: DPS: Our design pattern summarisation approach
@@ 5.1 RQ1: Statistical Evaluation of DPS against Human Summaries
Our study engaged a cohort of 25 participants, consisting of postgraduate computer science students and software engineering faculty members as mentioned, to generate human-written summaries of the corpus. We required each code sample to be summarised by a minimum of three independent participants to ensure reliability. While doing so, we asked participants to identify the design pattern they think the code contains and then summarise it. During the annotation and summarisation of the code, certain disagreements arose among participants regarding the classification of pattern types. For instance, in some cases, participants exhibited a tendency to identify a pattern as an abstract factory, whereas the correct classification was a factory method. Thus, to verify the level of agreement between these participants we calculated Cohen’s Kappa Carletta:1996 score, which was $>$ 0.7, i.e., there is moderate to high level of agreement between these participants.
@@
premaseem 0.38 7.2 0.48 0.51
Total 0.419 6.91 0.407 0.516
@@ 5.2 RQ2: Quality of Summaries
To further evaluate the quality of the DPS summaries, we asked human experts to rate the DPS generated summaries and human written summaries. We created a project on Prolific and shared a Google form that contained the source code, the summaries and five quality questions. The form contains a subset of selected summaries – at least one DPS summary and one human summary for each pattern – in total ten such instances. We asked the following five survey questions:
@@
The participants were given three options to choose from for each of the questions: 1) machine generated summary, 2) human written summary and 3) both. To minimise bias and maintain anonymity, the DPS generated summaries were labeled as Summary A, while the manually written summaries were labeled as Summary B. We set the following screening criteria for the participants: they must be computer science graduates, must be fluent in English, must have a postgraduate computer science degree or higher and at least 3 years of experience in using software design patterns in industrial settings. Around 15 participants responded to the survey and 10 were selected. The participants belonged to different demographics and fulfilled the screening criteria.
Based on the results, we observe that human-written summaries were rated higher in terms of accuracy, conciseness, and adequacy—reflecting the thoughtful effort and domain knowledge of our expert annotators. At the same time, DPS-generated summaries outperformed in two dimensions: capturing the context of the code and effectively conveying the underlying design pattern. These are precisely the aspects our approach aims to automate, and the results are encouraging. Despite being fully automated, DPS is able to match or even surpass human summaries in the structural and contextual dimensions that matter most for design pattern understanding. Table 5 shows the mean average values for each question for both summaries.
@@
Context captured 4.5 4
Convey DP 4 3
@@
6 Threats to the Validity

### v3

# ORCID: 0000-0003-2317-2297

@@ Abstract
Automatic summarisation has been used efficiently in recent years to condense texts, conversations, audio, code, and various other artefacts. A range of methods, from simple template-based summaries to complex machine learning techniques—and more recently, large language models—have been employed to generate these summaries. Summarising software design patterns is important because it helps developers quickly understand and reuse complex design concepts, thereby improving software maintainability and development efficiency. However, the generation of summaries for software design patterns has not yet been explored.
Our approach utilises code features and JavaParser to parse the code and create a JSON representation. Using an NLG library on this JSON representation, we convert it into natural language text that acts as a summary of the code, capturing the contextual information of the design pattern. Our empirical results indicate that the summaries generated by our approach capture the context in which patterns are applied in the codebase. Statistical evaluations demonstrate that our summaries closely align with human-written summaries, as evident from high values in the ROUGE-L, BLEU-4, NIST, and FrugalScore metrics. A follow-up survey further shows that DPS summaries were rated as capturing context better than human-generated summaries. Additionally, a time based task activity shows that summaries increase the time of understanding of design pattern for developer better than when the summaries are not present.
To our knowledge, this work presents the first automated approach to summarising software design patterns by linking pattern-detection features with natural-language generation. Unlike existing design-pattern detection or code summarisation methods that address these tasks separately, DPS introduces a feature-driven integration that converts detected pattern structures into contextual summaries describing the roles, relationships, and usage intent of each pattern instance.
@@ 1 Introduction
While previous research has either focused on detecting design patterns from source code or on generating summaries of general code elements, our work is the first to bridge these two directions. The novelty of DPS lies in (1) the construction of a feature-enhanced representation that encodes both class-level and method-level context for design-pattern instances, and (2) the generation of pattern-aware natural-language summaries that describe not only structural properties but also the contextual roles of participating classes. This integration enables automated documentation of design patterns, a problem that has not been previously studied.
@@
  • We introduce the first approach that integrates feature-based design-pattern detection with automatic natural-language generation to produce pattern-aware summaries describing the roles and relationships within design-pattern instances.
  • We construct a large corpus of 15 educational and industrial open-source Java systems (42,000+ files) containing implementations of nine GoF design patterns.
  • We extend existing feature-based detection techniques by expanding the feature set from 15 to 20 and using these features as structured inputs to a pattern-aware summarisation pipeline.
  • We empirically evaluate the approach using automatic metrics, expert judgement, and a comprehension study, showing that DPS summaries capture design-pattern context comparably or better than human-written summaries and improve developers’ comprehension efficiency.
The paper is organised into the following sections. Section 1 introduces the problem, and Section 2 provides the motivation behind the paper, particularly by discussing what context means. Section 3 discusses the experimental setup, focusing on corpus selection and generation, the code features used, the statistical measures for evaluation, and the research questions. Section 4 describes our three-phase summariser approach—DPS. Section 5 evaluates our work based on statistical measures and survey-based feedback. Section 6 discusses the results as well as the task based activity we performed. Section 7 discusses threats to internal and external validity, while Section 8 provides a description of related work in design pattern detection and code summarisation. Section 9 concludes our paper and discusses future directions for this research.
@@ 2 Background:Context
The notion of context in source code has been interpreted in various ways across the literature Mcburney:2014; Makharev:2025. In this work, we define context as the structural and behavioural information surrounding a code element (e.g., class or method) that is necessary to understand its role and functionality within a design-pattern instance. Specifically, we distinguish two types of context:
  • Method-level context, which captures the behaviour of a method through its dependencies, including the methods it invokes and the methods that invoke it.
  • Class-level context, which refers to the role a class plays within the pattern (e.g., acting as a Subscriber, Publisher, ConcreteFactory, or Product) and its structural relationships such as inheritance, composition, or interface implementation.
McBurney et al. Mcburney:2014 primarily focused on method-level context, defining it as the purpose of a method within the code in which it exists, including its dependencies and dependents. While this is essential for local comprehension tasks, in the case of design patterns the class-level context becomes even more critical, as it conveys the structural, behavioural, and creational intent of the pattern. For example, knowing whether a class functions as a ConcreteFactory or a Product in the Abstract Factory pattern strongly shapes how developers interpret its purpose and interactions.
Our DPS summarisation approach therefore captures both class-level and method-level context, representing them hierarchically: class-level roles and relationships form the high-level pattern narrative, while method-level behaviour is incorporated when it clarifies how these roles collaborate within the pattern.
The following example, taken from the RefactoringGuru GitHub repository, [^2] https://github.com/RefactoringGuru/design-patterns-java/blob/main/src/refactoring_guru/abstract_factory/example/factories/GUIFactory.java accessed and verified on 31-03-2025 illustrates the usage of context with respect to design patterns.
@@
The interface GUIFactory produces two products, namely Button and Checkbox. It’s implementation is shown below:
@@
In this example, the class-level context identifies GUIFactory as an abstract creator within the Abstract Factory pattern, while MacOSFactory serves as a concrete factory producing platform specific product variants. The method-level context explains how each method instantiates a particular product (MacOSButton, MacOSCheckbox). The structural context emphasises the interface (implementation) relationship that supports polymorphic behaviour and extensibility. Such contextual understanding is essential for tasks such as design pattern identification and comprehension.
Table 1 shows a same examples of design pattern contextual summaries taken from the DPS corpus.
@@
1 public class EmailNotificationListener implements EventListener { private String email; public EmailNotificationListener(String email) { this.email = email; } @Override public void update(String eventType, File file) { System.out.println("Email to " + email + ": Someone has performed " + eventType + " operation with the following file: " + file.getName()); } } EmailNotificationListener acts as a concrete observer inside the Observer Pattern, which updates String parameter of eventType, File parameter of file. It is a public class that implements EventListener. It sends email notifications. The only method of EmailNotificationListener is void update.
@@
4 public class HtmlDialog extends Dialog { @Override public Button createButton() { return new HtmlButton(); } } HtmlDialog is a part of factory method pattern. It is a public class that extends Dialog. It overrides the createButton() method to return an instance of HtmlButton. The only method of HtmlDialog is createButton (Button). createButton calls no methods.
5 public class BusAdapter implements Car { private Bus bus; public BusAdapter() { this.bus = new Bus(); } @Override public void drive() { bus.run(); } } BusAdapter acts as an Adapter, which overrides driveCar. It is a public class that implements Car. It adapts the Bus class to use a Car interface. BusAdapter implements the Car interface and internally holds a reference to a Bus object. The only method of BusAdapter is drive (void). No methods call drive and drive calls the run method.
@@ 3.1 Data Collection
To construct our corpus, we manually selected ten open-source educational repositories and five open-source general-purpose repositories from GitHub. The educational repositories were identified using the keyword “design patterns” and were specifically created to illustrate or teach software design patterns by the authors, rather than to serve as general-purpose software projects used by software developers. Many of these repositories originate from academic contexts (e.g., university course materials) or are associated with instructional resources such as books and websites—for example, the widely referenced RefactoringGuru repository.
While some educational repositories included examples for all 23 Gang of Four (GoF) design patterns, others featured a broader or narrower subset. We manually curated each repository to retain only those directories relevant to the nine design patterns targeted in our study. This involved selecting example code files from appropriate directories and excluding test cases, non-Java files, and unrelated content.
In contrast, the five general-purpose repositories—Eclipse JDT, Apache Hadoop, Spring Framework, Apache Camel, and jOOQ—are production-level projects that are actively maintained and used in industrial software development. They were selected to complement the educational examples with realistic, large-scale systems. We retained only those files that contained instances of the nine design patterns targeted in our study and excluded unrelated or non-Java files. The resulting dataset, referred to as the DPS-Corpus, now comprises slightly more than 42,000 Java files.
Table 2 presents the project names, their GitHub URLs, and the number of Java files containing implementations of the nine design patterns under investigation. The table summarises the distribution of design pattern instances across the fifteen Java projects included in our corpus. Each project demonstrates various design patterns, including Abstract Factory, Adapter, Decorator, Facade, Factory Method, Memento, Observer, Singleton, and Visitor. The total number of pattern-related files per project ranges from 46 to 6,079. Among the general-purpose repositories, Apache Camel contains the highest count while among the educational repositories, shihyu remains the largest and shusheng007 has the lowest count with 46 files. Patterns such as Abstract Factory and Factory Method are widely implemented across both educational and general-purpose repositories, whereas patterns like Decorator and Visitor appear less frequently.
@@
Spring Framework
spring-projects/spring-framework
114 81 23 510 146 20 26 76 19 1015
Apache Camel
apache/camel
1342 51 319 1830 1399 100 159 109 2 6079
Eclipse JDT
eclipse-jdt/eclipse.jdt.core
23 10 12 161 35 40 3 57 96 437
Apache Hadoop
apache/hadoop/
72 9 20 1294 100 282 67 451 38 2333
jOOQ
apache/hadoop
5 6 0 174 5 7 33 45 2 277
@@ 3.2 Code Features
Inspired by the work on Design Pattern Detection by Nazar et al. Nazar:2022, where they employ code features to identify design patterns, we have used code features to build the summaries. In this study, we have increased the number of features from 15 to 20 and renamed some of the features used by Nazar et al Nazar:2022. While some features (and their names) remain the same, others, such as NOML, are newly introduced. Additionally, features such as class name (CN) and method name (MN) have been renamed to enhance clarity and understanding. These features can be categorised into basic, object-oriented, and binary features, and further subdivided into class-level and method-level features. Feature extraction is a crucial aspect of code analysis, and accurately identifying a design pattern necessitates thorough code analysis. Table 3 lists the selected feature types, feature abbreviations, names, and descriptions.
@@
$O$ EXtendsFrom (EXF) List of classes that the Java file extends from, i.e. inherits
@@ 3.3.1 BLEU-4
$\text{BLEU-4}=\text{BP}\times\exp\left(\sum_{n=1}^{N}\tau_{n}\log P_{n}\right)$ (1)
@@ 3.3.2 NIST
$\text{NIST}=\sum_{n=1}^{N}\left\{\frac{\sum_{\text{all }n\text{-gram match that reference}}\text{Info}(n\text{-gram})}{\sum_{n\text{-gram reference}}}\right\}\cdot\exp(\beta)\log^{2}\left[\min\left(\frac{|\rho|}{|r|},1\right)\right]$
@@ 3.3.3 ROUGE-L
$\text{ROUGE-L}=\frac{(1+\beta)\times\text{Precision}\times\text{Recall}}{\text{Recall}+\beta^{2}\times\text{Precision}}$ (3)
@@ 3.3.4 FrugalScore
$\text{FrugalScore}=\alpha\cdot\text{EmbeddingSimilarity}+\beta\cdot\text{LexicalSimilarity}+\gamma\cdot\text{TaskSpecificComponent}$
@@ 4.1 Parsing
Figure 1: JSON representation of the ComputerPart class from AbdurRKhalid project (alt text: JSON excerpt showing keys for classes, methods, and pattern roles used by DPS to generate summaries)
@@ 4.2 Pattern Checker
It is important to note that DPS assumes the availability of a correctly detected design-pattern instance as input. The system operates downstream of established detection heuristics and focuses on summarising the structural and contextual features of these verified instances. While our pattern-checker module validates candidate instances using a 20-feature heuristic model, the goal of this component is not to propose a new detection technique but to ensure that summaries are generated from accurate pattern configurations. The summarisation quality therefore depends on the precision of the prior detection step.
@@ 4.3 Summary Generator
We used SimpleNLG, a Java-based natural language generation API, to produce design pattern summaries. This framework leverages lexical resources and realisation components to construct grammatically correct English sentences from structured inputs. The JSON output generated by our Pattern Checker module serves as the input for this phase, with SimpleNLG transforming the structured pattern information into coherent natural language summaries.
@@
Figure 2: DPS: Our design pattern summarisation approach (alt text: block diagram showing parsing to AST, pattern checker producing JSON, and summary generator producing natural-language descriptions)
@@ 5.1 RQ1: Statistical Evaluation of DPS against Human Summaries
Our study engaged a cohort of 25 participants, consisting of postgraduate computer science students and software engineering faculty members as mentioned, to generate human-written summaries of the corpus. We required each code sample to be summarised by a minimum of three independent participants to ensure reliability. While doing so, we asked participants to identify the design pattern they think the code contains and then summarise it. During the annotation and summarisation of the code, certain disagreements arose among participants regarding the classification of pattern types. For instance, in some cases, participants exhibited a tendency to identify a pattern as an abstract factory, whereas the correct classification was a factory method. Thus, to verify the level of agreement between these participants we calculated Cohen’s Kappa Carletta:1996 score, which was >> 0.7, i.e., there is moderate to high level of agreement between these participants.
@@
premaseem 0.42 7.2 0.48 0.51
Spring Framework 0.64 7.6 0.66 0.62
Apache Camel 0.62 6.9 0.67 0.63
Apache Hadoop 0.59 7.4 0.58 0.6
Eclipse.jdt.core 0.65 7.8 0.64 0.68
jOOQ 0.58 8.1 0.71 0.61
Total (avg) 0.5 7.17 0.51 0.55
@@ 5.2 RQ2: Quality of Summaries
To further evaluate the quality of the DPS summaries, we asked human experts to rate the DPS generated summaries and human written summaries. We set the following screening criteria for the participants: they must be computer science graduates, must be fluent in English, must have a postgraduate computer science degree or higher and at least 3 years of experience in using software design patterns in industrial settings. Around 20 participants responded to the survey and 15 satisfied our criteria. The participants belonged to different demographics and fulfilled the screening criteria.
To mitigate potential bias, all survey participants were anonymised and were not informed which summary was human-written or automatically generated. Each code example was presented with two unlabeled summaries—Summary A and Summary B—and participants answered five quality questions without knowing the source of either summary. Participants were screened to ensure at least three years of professional experience using design patterns in industrial contexts, postgraduate-level computer-science education, and Java proficiency.
We asked the following five survey questions:
@@
The participants were given three options to choose from for each of the questions: 1) machine generated summary, 2) human written summary and 3) both.
Based on the results, we observe that human-written summaries were rated higher in terms of accuracy, conciseness, and adequacy—reflecting the thoughtful effort and domain knowledge of our expert annotators. At the same time, DPS-generated summaries outperformed in two dimensions: capturing the context of the code and effectively conveying the underlying design pattern. These are precisely the aspects our approach aims to automate, and the results are encouraging. Despite being fully automated, DPS is able to match or even surpass human summaries in the structural and contextual dimensions that matter most for design pattern understanding.
Table 5 reports mean scores averaged across all participants for each evaluation dimension (accuracy, conciseness, adequacy, context captured, and pattern conveyed). These values represent the aggregated perception of quality rather than direct pairwise votes.
@@
Context captured 5.5 4
Convey DP 5 3
@@
6 Discussion
As shown in Table 5, the proposed Design Pattern Summaries (DPS) achieve accuracy on par with human summaries (4.5 each) while surpassing them in capturing contextual information (5.5 vs.4) and conveying the design pattern itself (5 vs.3), even though human summaries score higher in conciseness and adequacy.
To illustrate these differences more concretely, we provide six representative examples of summary outputs: three instances where DPS-generated summaries outperform human-written ones, and three instances where human-generated summaries are judged to be more effective than those produced by DPS.
Table 6 shows three instances of DPS summaries performed better than human summaries.
Table 6: Three instances of summaries where DPS performed better than Human Summaries.
URL DPS Summary Human Summary
EmailNotificationListener acts as a concrete observer inside the Observer Pattern, which updates String parameter of eventType, File parameter of file. It is a public class that implements EventListener. It sends email notifications. The only method of EmailNotificationListener is void update. The EmailNotificationListener is a concrete observer as it receives notifications about events related to file operations. It overrides the update method that sends email notifications.
The Dialog class acts as an abstract creator in the factory method pattern. Dialog acts as a factory method for Button, which is inherited by Dialog, It is a public abstract class. The 2 methods of Dialog are renderWindow (void) and createButton (Button). The createButton is overridden by all subclasses for button types for each type of dialog. No methods call renderWindow and renderWindow calls only one method : createButton method of class Dialog. Only one method (renderWindow method of class Dialog) calls createButton and createButton calls no internal methods. The Dialog class is an Abstract Creator providing a template for creating dialogs that can generate buttons.
DefaultINodeAttributesProvider acts as a caretaker for memento Attribute, which starts, which stops, which gets String[] parameter of pathElements, INodeAttributes parameter of inode Attributes. It has three methods start(), stop() and getAttributes(). It calls no method. The DefaultINodeAttributesProvider class in Apache Hadoop HDFS provides a default implementation of INodeAttributeProvider. It includes empty start and stop methods and a getAttributes method that returns the provided INodeAttributes without modification, serving as a basic, no-operation attribute provider for HDFS inodes.
textcolorblueTable 7 shows examples where human summaries were better than DPS summaries.
Table 7: Three instances of summaries where DPS underperformed than Human Summaries.
URL
DPS Summary Human Summary
Computer.java Computer does not have any design pattern. It is a public class that extends ComputerPart and that implements ComputerPart. The only method of Computer is accept (void). No methods call accept and accept calls no methods. The Computer class represents a Concrete Element of Visitor, which allows operations to be performed on Mouse, Keyboard and monitor components by an external visitor.
ScopedProxyBeanDefinitionDecorator does not follow any design pattern. It is a package-private class that implements BeanDefinitionDecorator. The only method is decorate(Node, BeanDefinitionHolder, ParserContext). No methods call decorate and decorate calls a utility to make a proxy. The ScopedProxyBeanDefinitionDecorator class applies the Decorator design pattern within the Spring Framework’s AOP configuration context. It customises bean definition parsing by intercepting the ¡aop:scoped-proxy/¿ tag and wrapping the original bean with a scoped proxy. This enables context-aware proxying of Spring beans, particularly useful for injecting request or session-scoped beans into singleton scoped components while preserving AOP capabilities.
Member does not have any design pattern. It is a public interface. The only method of Member is getDescription (String). No methods call getDescription and getDescription calls no methods. The Member interface defines a contract for ship members or related entities, requiring them to provide a description via the getDescription() method, supporting the Abstract Factory pattern’s consistency.
The quantitative results in Table 4 reinforce that across 15 diverse projects, DPS achieved average scores of 0.50 for BLEU‑4, 7.17 for NIST, 0.51 for ROUGE‑L, and 0.55 for FrugalScore, indicating strong alignment with reference summaries. High NIST values for projects such as Luisburgos (8.5) and shihyu (8.3) highlight the method’s ability to capture informative n‑grams, while elevated ROUGE‑L scores for jOOQ (0.71) and Apache Camel (0.67) reflect substantial overlap in longer, semantically meaningful sequences. Altogether these results demonstrate that the proposed approach not only produces summaries competitive with human‑authored ones but also excels in preserving the contextual integrity of design patterns.
To complement the subjective survey, we performed a task-based comprehension study to evaluate whether automatically generated summaries assist in understanding design-pattern roles within source code. A research engineer with three years of professional experience was asked to identify all participating roles for nine design patterns, both with and without summaries. Each pattern instance was selected randomly to avoid order effects. Across all patterns, the participant completed the tasks more quickly when summaries were provided (see Table 8), while accuracy remained constant. This finding demonstrates that DPS summaries improve comprehension efficiency for realistic program-understanding tasks.
For each design pattern, we randomly selected ten instances from the dataset. In each instance, the participant was provided with a GitHub link to a single class known to play one role in that pattern (e.g., the AbstractFactory class). The participant’s task was to identify all other classes corresponding to the remaining roles of that pattern. For example, for the Abstract Factory pattern, the participant had to find the AbstractFactory, ConcreteFactory, AbstractProduct, and ConcreteProduct roles.
The study followed a within-subject design: for half of the instances (five per pattern), the participant was shown our automatically generated summary; for the other half, no summary was provided. All assignments were randomly selected to avoid order or learning effects. The participant was allowed to consult the Refactoring Guru definition (https://refactoring.guru/design-patterns) of each pattern during the task. Time to completion was self-recorded for each instance. Each instance was attempted once, and no outliers were excluded.
We excluded any cases where fewer than three roles were present. Table 1 shows the average time (in minutes and seconds) required to complete the role-identification task, comparing the conditions with and without summaries.
Table 8: Role Identification Summary results with and without summaries
Design Pattern Without Summary With Summary
Abstract Factory 2:26 1:39
Adapter 3:26 2:59
Factory Method 3:01 0.59
Visitor 2:23 1:48
Across all four patterns, tasks performed with summaries were completed more quickly than those without. Accuracy was consistent across conditions, indicating that the improvement was due to efficiency rather than guesswork or omission. These results suggest that the summaries helped the participant locate and recognize pattern roles more efficiently during program comprehension.
Large language models (LLM) based tools such as ChatGPT have shown considerable potential in natural language generation; however, their performance in producing design pattern contextual summaries is inherently sensitive to prompt engineering. Prompt engineering introduces a substantial experimental dimension, requiring systematic design, tuning, and evaluation to ensure fairness and reproducibility. Incorporating such models would therefore expand the scope well beyond the objectives of this study. Our focus is instead on a domain specific, code feature driven approach that can be applied consistently without the variability introduced by prompt engineering dependent systems. This enables a more controlled evaluation of summary quality.
7 Threats to Validity